HUBBLE SPACE TELESCOPE DAILY REPORT #1591 PERIOD COVERED: 0000Z (UTC) 03/14/96 - 0000Z (UTC) 03/15/96 Daily Status Report as of 075/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.2 Completed GHRS/1 5891 (Structurally Resolved Abundances and = Depletions in the Rho Oph Cloud) The High Resolution Spectrograph (Side One) was used to make high = resolution spectral observations of star HD143275 which is located within = the Rho Orpiuchi Cloud in order to study its structurally resolved = elemental abundances and depletions in different parts of the cloud. The = observations were completed as scheduled, and no anomalies were reported. 1.3 Completed Two Sets of FOS/RD 5823 (Uranus: Ringlet Kinematics and = Atmospheric Structures) The Faint Object Spectrograph (Red Side) was used to make high = resolution spectrographic observations of the planet Uranus, its moon = Titania, and a ringlet segment as a background star passed through the = Uranian System. The observations will provide information about the = planet's atmosphere and its rings. The observations were executed as = scheduled, and no anomalies were noted. 1.4 Completed GHRS/1 5882 (Ech-A Studies of Single Interstellar = Clouds) The High Resolution Spectrograph (Side One) was used to make = Echelle-A spectrographic observations of the star HD141637 (aka, Rho Leo) = in order to study the interstellar absorption line clouds which lies in = the line of sight to that star. The observations were completed as = planned, and no problems were reported. 1.5 Completed WF/PC-2 5913 (Low-Mass Binaries and Luminosity Fuction) The WF/PC-2 was used to make observations of the low-mass binary = system NGP016 (Late Type Class M dwarf stars). The intent is to = determine why the local stellar luminosity function is flat below = Mv=3D12, while that found from more distant stars falls by an order of = magnitude by Mv=3D14. The observations were executed as scheduled, and = no anomalies were noted. 1.6 Completed COSTAR/U 6162 (FOC Cycle 5 PSF Monitoring) The Faint Object Camera (f/96) was used to make observations of a = standard star and to monitor the point spread function of the COSTAR = optical system. The observations were completed as scheduled, and no = anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 6363 (The Environments and Host = Galaxies of BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of = BL-Lac object S5-054+84 in order to study its environment and its host = galaxy. The goal is to investigate systematic trends in host galaxy = types, unusual morphological features, nearby companion galaxies, etc. = The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed GHRS/2 6086 (Searching For Flaring Activity on the Star = Gamma Cas) The High Resolution Spectrograph (Side Two) was used to make rapid = read-out mode spectrographic observations of the Be class star HD5394 = (aka, Gamma Cas) to search for flaring activity. This star shows = variability in x-rays. A wavelength calibration image was also taken. = The observations were completed as planned, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 5 Successful : 5 2.3 Operations Notes: The secondary mirror was moved +6 microns to correct for metering = truss desorption at 074/1848Z. The next move is anticipated in = approximately 7 months. The STR dump commanded at 075/0505Z was expected to contain both = science and engineering data. The engineering data was not present due = to inconsistencies between the 32K record in the SMS and a 4K format = commanded in real time for the mirror move. There was a 5 minute 2 second non-recoverable loss of data to a = low rate switch failure at STGT. Later, there was a 13 minute 11 second = recoverable 1M data loss due to a NASCOM data base problem. TTRs were = written for both cases. 3.0 SIGNIFICANT FORTHCOMING EVENTS: NSSC-1 software patch 5.3B is expected to be installed at = approximately 075/1957Z. /DMH